# Many-body density of states of bosonic and fermionic gases: a combinatorial approach

### v1

@@ 1 Introduction
Recent experiment with cold atoms [1, 2] and the interest in many-body (MB) systems that involve high excitations, such as MB thermalization [3, 4, 5] and MB localization [6, 7], have rekindled the need for analytical estimates of the many-body density of states (MBDOS) and its fluctuations [8, 9, 10]. In particular, the celebrated Bethe estimate for the mean level density of fermionic systems in the mean-field regime has drawn much recent attention due to its ubiquitous presence in systems with a holographic gravitational dual like SYK [11] and also in low-dimensional dilaton gravity [12].
@@
The paper is organized as follows. Sec. 2 outlines the derivation and the range of validity of the Bethe approximation and of its bosonic analogue. Sec. 3 is dedicated to discussing the combinatorial solution to the MBDOS in the case of equally spaced SP spectra. Eq. (49) summarizes the main result, and its derivation through Eq. (47) closes a gap in the existing literature. We show that the Bethe formula and its bosonic counterpart are recovered from the combinatorial results in a limit consistent with their validity range. Further limits are discussed. The conclusions are presented in Sec. 4.
@@ 2.1 Standard derivation of the MBDOS of a fermionic gas
where $g(\epsilon)\sum_{\nu}\delta\!\left(\epsilon-\epsilon_{\nu}\right)$ is the system SP level density. Assuming that the interval in which the integrand of Eq. (6) is notably different from zero is wide compared to the spacing of the SP levels, one may replace the SP level density by a smooth function of the energy, here also called $g$. If at the same time, this window is narrow compared to the scale over which $g$ varies, the problem satisfies the conditions of the Sommerfeld integral [21]. Hence, $\text{log}~{}Z(\alpha,\beta)$ can be written as
@@ 3 MBDOS of systems with constant SPDOS
Quantum indistinguishability adds a layer of complexity.
@@
In the case of fermionic particles, the Pauli exclusion principle imposes the additional constraint that the $m_{i}$ be distinct. In this case, the combinatorial problem is solved by the number of partitions of ${\cal E}$ into exactly $N$ or $N-1$ distinct positive integer parts
$d_{=N}({\cal E})+d_{=N-1}({\cal E})$ (42)
counting the number of multisets with distinct entries of which none or precisely one is zero, corresponding to whether or not the ground level is occupied. This is illustrated in Fig. 1.
Figure 1: Fermionic (a) and bosonic (b) MB configurations of $N$ harmonic oscillators at excitation energy $\hbar\omega{\cal Q}$, here $N=3$ and ${\cal Q}=5$. Each pannel represents, within the occupation number picture, a state of the correct symmetry with SP energies $\left(m_{i}+\frac{1}{2}\right)\hbar\omega$, $1\leq i\leq 3$, indicated by filled circles. The vertical arrows display the SP excitations with respect to the system ground state, indicated by dashed circles. They illustrate the one-to-one mapping between fermionic and bosonic configurations. The horizontal blue dashed line (in (a)) stands for the Fermi energy level.
@@ Acknowledgments
We thank Quirin Hummel for useful conversations. CL thanks the Brazilian funding agencies CNPq and FAPERJ. GM is funded through a fellowship by the Studienstiftung des Deutschen Volkes. We acknowledge financial support from the Deutsche Forschungsgemeinschaft (German Research Foundation) through Ri681/15-1 (project number 456449460) within the Reinhart-Koselleck Programme.

### v2

@@ 1 Introduction
Recent experiments with cold atoms [1, 2] and the interest in many-body (MB) systems that involve high excitations, such as MB thermalization [3, 4, 5] and MB localization [6, 7], have rekindled the need for analytical estimates of the many-body density of states (MBDOS) and its fluctuations [8, 9, 10]. In particular, the celebrated Bethe estimate for the mean level density of fermionic systems in the mean-field regime has drawn much recent attention due to its ubiquitous presence in systems with a holographic gravitational dual like SYK [11] and also in low-dimensional dilaton gravity [12].
@@
The paper is organized as follows. Sec. 2 outlines the derivation and the range of validity of the Bethe approximation and of its bosonic analogue. Sec. 3 is dedicated to discussing the combinatorial solution to the MBDOS in the case of equally spaced SP spectra. Eq. (49) summarizes the main result, and its derivation through Eq. (47) closes a gap in the existing literature. We show that the Bethe formula and its bosonic counterpart are recovered from the combinatorial results in a limit consistent with their validity range. Further limits as well as prospects of generalization are discussed. The conclusions are presented in Sec. 4.
@@ 2.1 Standard derivation of the MBDOS of a fermionic gas
where $g(\epsilon)=\sum_{\nu}\delta\!\left(\epsilon-\epsilon_{\nu}\right)$ is the system SP level density. Assuming that the interval in which the integrand of Eq. (6) is notably different from zero is wide compared to the spacing of the SP levels, one may replace the SP level density by a smooth function of the energy, here also called $g$. If at the same time, this window is narrow compared to the scale over which $g$ varies, the problem satisfies the conditions of the Sommerfeld integral [21]. Hence, $\text{log}~{}Z(\alpha,\beta)$ can be written as
@@ 3 MBDOS of systems with constant SPDOS
Quantum indistinguishability adds a layer of complexity:
@@
In the case of fermionic particles, the Pauli exclusion principle imposes the additional constraint that the $m_{i}$ be distinct. In this case, the combinatorial problem is solved by the number of partitions of ${\cal E}$ into exactly $N$ or $N-1$ distinct positive integer parts, denoted $d_{=N}({\cal E})$ and $d_{=N-1}({\cal E})$ respectively, hence the sum
$d_{=N}({\cal E})+d_{=N-1}({\cal E}).$ (42)
Eq. (42) counts the number of multisets with distinct entries of which none or precisely one is zero, corresponding to whether or not the ground level of the physical system is occupied. This is illustrated in Fig. 1 for the specific case of $N=3$ and an excitation energy $Q=\hbar\omega{\cal Q}$ with ${\cal Q}=5$.
Figure 1: Fermionic (a) and bosonic (b) MB configurations of $N$ harmonic oscillators at excitation energy $\hbar\omega{\cal Q}$, here $N=3$ and ${\cal Q}=5$. Each pannel represents, within the occupation number picture, a state of the correct symmetry with SP energies $\left(m_{i}+\frac{1}{2}\right)\hbar\omega$, $1\leq i\leq 3$, indicated by filled circles, and subject to the constraint (a) $\sum_{i=1}^{3}m_{i}=5+(0+1+2)=8$, (b) $\sum_{i=1}^{3}m_{i}=5$. The number of independent admissible states is seen to be given by (a) $d_{=3}(8)+d_{=2}(8)=2+3=5$, (b) $p_{\leq 3}(5)=5$. The horizontal blue dashed line (in (a)) stands for the Fermi energy level. The vertical arrows display the SP excitations with respect to the system ground state, indicated by dashed circles. They illustrate the one-to-one mapping between fermionic and bosonic configurations, thereby demonstrating that the MB level degeneracies are the same for the two particle species.
@@ 3.3 Distinguishable particles and the correct Boltzmann counting
3.4 Prospects of generalization
While the discussion at the beginning of the present section, leading to Eq. (37), applies to an arbitrary non-interacting MB system, determining the MB level degeneracies is, in general, a difficult combinatorial problem. The case of constant SPDOS considered in this work has the virtue that the counting problem becomes tractable, and a natural question is to what extent the combinatorial approach and its results can be extended to other physical systems.
For $N$ particles under harmonic confinement in more than one, say $D$, spatial dimensions, computing the degeneracies remains a linear integer partition problem, admitting a similar analysis as conducted here. For distinguishable particles, the system is equivalently described as an assembly of $ND$ one-dimensional harmonic oscillators and therefore covered by our results Eq. (49), whereas for indistinguishable particles, modified constraints on the parts of the integer partition have to be taken into account.
An energy-dependent mean SP level spacing, in contrast, adds considerably to the complexity of the counting problem. Consequences of fluctuations about a rigid spectrum are investigated in Refs. [13, 36]. For SP energies obeying a power law in the quantum number, the problem of computing the MB level degeneracies translates into a non-linear partition problem, that is the counting of partitions into powers of integers, for which asymptotic results are obtained in Ref. [37]. We remark that the finding to which the results of Ref. [37] specialize in the case of the harmonic oscillator and which in our notation reads $d_{=N}\!\left({\cal Q}+\frac{N^{2}}{2}\right)\sim p_{\leq N}({\cal Q})$ for large $N$ and ${\cal Q}$, can also be deduced from Eq. (62).
@@ Acknowledgments
We thank Quirin Hummel for useful conversations. CL thanks the Brazilian funding agencies CNPq and FAPERJ. GM is funded through a fellowship by the Studienstiftung des Deutschen Volkes. We acknowledge financial support from the Deutsche Forschungsgemeinschaft (German Research Foundation) through Ri681/15-1 (project number 456449460) within the Reinhart-Koselleck Programme. Finally, we would like to thank the anonymous referees for their comments, which led to improvements of the manuscript and prompted us to include subsection 3.4.